OpenStudy (anonymous):

(Can someone help me understand this problem?) Andrea has a yard shaped like parallelogram ABCD. The garden area in the corner is represented by parallelogram EFGB. Andrea wants to put a fence around her yard. She also wants to fence in the garden. How many feet of fence will she need? A.159 ft B.189 ft C.210 ft D.216 ft

OpenStudy (anonymous):



ganeshie8 (ganeshie8):

find the length around the yard ABCD

OpenStudy (anonymous):

i started out with 12^2+15^2= 369. Then i got lost..

OpenStudy (anonymous):

45+45+36+36= Like this?

ganeshie8 (ganeshie8):

yes, wat do u get

OpenStudy (anonymous):

90+72=162

ganeshie8 (ganeshie8):

yes thats the length of fence around outer yard

ganeshie8 (ganeshie8):

ne need to fence garden also eh ?

OpenStudy (anonymous):

wait what??.. am i missing a number?

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

just add that length also to 162

ganeshie8 (ganeshie8):

total fence = 162 + 12 + 15

OpenStudy (anonymous):

ohh... 189

ganeshie8 (ganeshie8):

Yes !

OpenStudy (anonymous):

no wonder i kept getting it wrong! I was squareing. Thanks ganesh! :D

ganeshie8 (ganeshie8):

np :)